OpenStudy (anonymous):

A room is elliptical. If it is 20 feet long (horizontally) and 18 feet wide (vertically), how far apart would two people be if they are standing at the foci?

OpenStudy (amistre64):

20^2 = 18^2 + foci^2 solve for foci and double it :)

OpenStudy (amistre64):

oh.... and use the right numbers lol; 10 and 9

OpenStudy (amistre64):

i get 8.718 abt :)

OpenStudy (amistre64):

\[2\sqrt{(20/2)^2 - (18/2)^2}=distance\]